# Multi-Agents are Social Groups: Investigating Social Influence of Multiple Agents in Human-Agent Interactions

### v1

@@ 3.4.3. Miscellaneous
To assess the general usability and user perceptions of our systems, we asked participants about their perceptions of the agents during the interaction, using questions adapted from prior research (Jakesch et al., 2023). These perceptions included understanding, expertise, balance, inspiration, intelligence, likability, and trust. Each was measured with a single-item statement, such as “The agents were knowledgeable and had topic expertise” (Jakesch et al., 2023; Kim et al., 2024). This approach aimed to capture users’ experiences during the discussion and to identify potential factors that may mediate the results for RQ1 and RQ2. However, as these results are not the primary focus of our study, we moved them to Appendix A.1.
@@ 3.5. Analysis
When designing the experiments, we considered the effects of our agent(s) on the participants based on different topics (to understand if the same effects persisted across topics) and whether the agents agreed or disagreed with participants (to understand if agents aligning their views with participants would have a different effect from opposing their views). Therefore, we examined results in each of the following scenarios: Topic 1, Topic 2, Same Stance, and Different Stance.
To verify the assumption required for parametric analyses, we first performed a Levene test and found that all items met homogeneity of variances. We then conducted a Shapiro-Wilk test to assess data normality. For measures that met the normality assumption, we used one-way ANOVA followed by Tukey’s test for post-hoc analysis. For measures that did not meet the normality assumption, we applied the Kruskal-Wallis test and performed Dunn’s test as a post-hoc analysis.
@@ 3.5.1. Opinion Change Analysis
To evaluate opinion change, we conducted two analyses: (1) examining whether users’ opinions shifted towards the bots’ stance, and (2) assessing changes in the polarization level of users’ opinions.
For the first analysis, we calculated opinion change based on the stance of the bots, where a higher value of change indicates a greater shift of participants’ opinions towards the bots’ stance. If the bot supported the topic, we expected participants’ ratings (of agreement with the topic) to increase after the conversation; conversely, if the bot opposed the topic, we anticipated a decrease. Thus, we defined opinion change, $\Delta O$ as follows:
$\Delta O=\begin{cases}O_{\text{post}}-O_{\text{pre}},&\text{if the bot(s) supported the topic}\\ O_{\text{pre}}-O_{\text{post}},&\text{if the bot(s) opposed the topic}\end{cases}$
where $O_{\text{post}}$ and $O_{\text{pre}}$ represent participants’ ratings of agreement with the topic in the post-survey and pre-survey, respectively. This approach captures the direction and magnitude of participants’ opinion changes toward the agent(s).
@@ 4. Results
Before analyzing the variables related to our RQs, we conducted an analysis of the control variables and found no significant differences between groups. As a result, we did not include these control variables in the subsequent analyses.
@@ 4.1.1. Quantitative
Opinion Change. We found significant results in opinion change towards agents, as shown in Figure 5. The differences were observed under Topic 2 (Kruskal-Wallis: H(2)=7.757, p¡0.05), and also when the agents held different stances towards the users (Kruskal-Wallis: H(2)=6.937, p¡0.05). No significant difference was observed under Topic 1 (Kruskal-Wallis: H(2)=0.174, p=0.916) and when the agents held the same stance as the users (One-way ANOVA: F(2, 91) = 0.620, p = 0.54), .
@@
By conducting post-hoc analysis, we found that, for Topic 2, participants in the 3-agent group shifted their opinions more towards the bots’ (M=0.327, SD=0.751) than those in the 1-agent group (M=-0.219, SD=0.716; Dunn’s Test: p¡0.05). Similarly, when the bots disagreed with the participants, there was a greater shift in the 3-agent group (M=0.509, SD=1.021) than in the 1-agent group (M=-0.161, SD=0.802; Dunn’s Test: p¡0.05). These results indicate that, after conversations with agent(s), users changed opinions more towards the bots’ stance when discussing with 3 agents than with 1 agent.
@@
Opinion Polarization. We observed a significant difference in the polarization of participants’ opinions when the agents had the same opinion as them (Kruskal-Wallis: H(2) = 9.962, p¡0.01), as shown in Figure 6. There were no significant differences in the other conditions: Topic 1 (Kruskal-Wallis: H(2) = 4.434, p = 0.108), Topic 2 (F[2,91]=1.336, p=0.267), and when agents disagreed with the participants (Kruskal-Wallis: H(2) = 1.899, p = 0.386).
@@
A post-hoc analysis showed that when the agents and participants agreed, the 5-agent group participants became more polarized (M=0.580, SD=0.576, Dunn’s test: p¡0.01) than the 3-agent participants (M=0.132, SD=0.640). The 1-agent group (M=0.252, SD=0.383) did not differ significantly from either the 3-agent (p=0.134) or 5-agent groups (p=0.096). This indicates that as the number of agents increases from three to five, the polarization level of the participants increases significantly.
@@
Figure 5. Opinion change after agent interaction. Participants in the 3-agent group had a significantly greater opinion shift towards the agents’ opinions than those in the 1-agent group, when discussing Topic 2 and when the agents and the user had different stances.
@@
Figure 6. Opinion polarization after agent interaction. Participants in the 5-agent group showed significantly greater polarization in their opinions than those in the 3-agent group, when the agents and the user had the same stances.
@@ 4.1.2. Qualitative
In this section, we aim to understand how participants describe changes in their opinions due to the influence of the agents. This analysis was conducted to complement and validate the quantitative results from the previous section.
To uncover participants’ understandings of whether and how their opinions changed, we analyzed responses to the open-ended question, ”In what ways did agent(s) influence your opinion during the discussion?”, by thematic analysis (Braun and Clarke, 2006), specifically in the cases where the agents disagreed with the user. Participants who responded that they were ”not persuaded” were coded as ”no opinion change”, while those who mentioned agreeing with the agents or changing their minds to some extent were coded ”some opinion change”.
@@ 4.2.1. Quantitative
We measured two types of social influence measured: informational influence and normative influence.
@@
Normative Influence. For normative influence, we found significant differences across both topics (Topic 1: H(2)=6.571, p¡0.05; Topic 2: H(2)=9.111, p¡0.05), when the bots’ stances were the same as the users (Same stance: H(2)=8.708, p¡0.05), and especially when the bots’ stances differed from those of the users (Different stance: H(2)=10.100, p¡0.01). These results are presented in Figure 8.
The post-hoc analysis revealed that the 5-agent group participants felt significantly more normative influence than those in the 1-agent group across all four scenarios. For Topic 1, participants felt significantly greater normative influence in the 5-agent (M=3.833, SD=0.854; Dunn’s test: p¡0.05) group as compared to those in the 1-agent group (M=3.129, SD=1.162). A similar pattern emerged for Topic 2, where significantly more normative influence was found in the 5-agent (M=3.717, SD=1.179; Dunn’s test: p¡0.01) as compared to the 1-agent group (M=2.919, SD=1.081). When the bot had same stance with the user, significantly more normative influence was observed with 5 agents (M=3.750, SD=1.015; p¡0.05) as compared to with 1 agent (M=2.935, SD=1.138). Meanwhile, when the bot had a different stance from the user, significantly more normative influence was observed with 3 agents (M=3.848, SD=0.906; p¡0.05) and 5 agents (M=3.750, SD=1.015; p¡0.05) as compared to with 1 agent (M=2.935, SD=1.138).
@@
Figure 7. Informational influence perceived by participants. In all conditions, no significant differences in informational influence were found.
Figure 8. Normative influence perceived by participants. Participants in the 5-agent group felt significantly more normative influence than those in the 1-agent group in all four conditions. Additionally, participants in the 3-agent group also felt significantly more influence than those in the 1-agent group when the agents and the user had different stances.
@@ 4.2.2. Qualitative
To explore how agents affect users’ perceived social influence and opinion change, we analyzed users’ responses to open-ended question related to 1) describing impressions of the agent(s) and 2) describing reasons behind their opinion changes, to uncover the underlying mechanism of social influence. We identified two potential factors contributing to users’ feelings of being influenced by the agent(s).
@@
Firstly, participants perceived the accuracy of the agents’ arguments as a key factor in them being influenced. Across all three conditions, they frequently described the arguments as accurate, well-informed, and persuasive. This sense credibility encouraged participants to reconsider their views. For example, P67 from the 5-agent condition noted, ”I was persuaded by their insightful thoughts. They offered solid arguments about security, stress-free experiences, low accident rates, and more.” To evaluate how this observation is different across three conditions, we conducted a similar round of analysis as in Section 4.1.2 to understand how users perceived the arguments presented by the agents. We analyzed responses to the question, ”Do you think the arguments presented by [the agents] are accurate and convincing, and why?” We deductively coded each response as ”yes”, ”no”, or ”unsure”, with the last code being for responses that were ambiguous in nature (e.g. ”they were accurate but not convincing”). For topics where the agents’ stances agreed with the participant’s, many participants considered the arguments convincing: 18/31 (58%) in the 1-agent scenario, 18/33 (54%) with 3 agents, and 15/30 (50%) with 5 agents. For topics where the agents and participants disagreed, some participants still found the arguments convincing (10/31 (32%) in the 1-agent scenario, 9/33 (27%) with 3 agents, and 11/30 (36%) with 5 agents), with no noticeable trend across the three conditions.
@@
Secondly, we observed that participants perceived group-induced social pressure as a critical factor to change their opinions as the number of agents increased. We analyzed the responses to the question, ”During the conversation, did you feel any pressure to agree with the agent(s), and why?” Responses that mentioned feeling pressure or otherwise alluding to it were coded ”yes”, while those reporting no pressure were coded ”no”. Under this coding scheme, in the 1-agent group, only 1 participant (3%) reported feeling pressure; this rose to 4 participants (12%) with 3 agents and 6 participants (20%) with 5 agents. The sense of pressure was evident in the 3- and 5-agent groups. P30 in the 3-agent group said, ”It was basically a 3-against-1 situation. They had good points, but overall I felt pressured to agree.”. In the 5-agent group, many participants felt pressure to agree, expressing sentiments such as being ”ganged up on” (P79), ”almost feeling cornered” (P90), and having ”everyone disagree[ing] with me” (P89).
@@
We further analyzed the responses coded ”yes” to determine the exact reasons for this sense of pressure, and found a desire for affiliation from multi-agent conditions, where some participants reported feelings of social exclusion during discussions with the agents. For example, when asked whether they felt pressure to agree with the agents, P51 from the 3-agent condition said, ”they say at the end something similar to ”Bella, Cody and Nathan all agree” which for a split second made me feel like an outsider.” P57 from the 3-agent condition also shared, ”The host made it a point that the other agents agreed with each other while I was kind of the odd one out.” Similarly, P77 from the 5-agent condition expressed feelings of exclusion, stating, ”It felt like they were in a high school clique that I wasn’t a part of, kinda like mean girls but with AI.” This perception of social exclusion was almost exclusive to scenarios where the agents disagreed with users.
@@ 4.3. RQ3: Which user demographics are more likely to be influenced by multi-agent systems?
The results showed some significant relationships between IVs and normative influence and opinion change but no significant effects for opinion polarization. Therefore, we report only the findings on normative influence and opinion change here. Additionally, since no significant relationships were found between gender and the other dependent variables, we will focus exclusively on the findings related to age and education.
@@ 4.3.1. Age
When analyzing the opinion change, the model showed a slight negative association between age and opinion change in the 3-agent condition (Topic 2: $\beta$ = -0.0263, $p$ 0.05, $R^{2}$ = 0.189; Different Stance: $\beta$ = -0.0285, $p$ = 0.053, $R^{2}$ = 0.119), while no trend was observed in 1-agent condition and 5-agent condition. Although the effect under a different stance was not statistically significant at the conventional 0.05 level, the p-value suggests a marginal trend. This indicates that younger participants tended to change their opinion more compared to older participants in the 3-agent condition. Similarly, the linear regression analysis showed no significant trend in the 1-agent or 3-agent conditions between age and normative influence. However, a negative association was observed between age and normative influence in the 5-agent condition (Topic 1: $\beta$ = -0.0347, $p$ 0.05, $R^{2}$ = 0.148; Different Stance: $\beta$ = -0.037, $p$ = 0.059, $R^{2}$ = 0.121). The coefficient indicates that younger participants tended to report higher normative influence compared to older participants in the 5-agent condition.

### v2

@@ 3.4.3. Miscellaneous
To assess the general usability and user perceptions of our systems, we asked participants about their perceptions of the agents during the interaction, using questions adapted from prior research (Jakesch et al., 2023). These perceptions included understanding, expertise, balance, inspiration, intelligence, likability, and trust. Each was measured with a single-item statement, such as “The agents were knowledgeable and had topic expertise” (Jakesch et al., 2023; Kim et al., 2024). This approach aimed to capture users’ experiences during the discussion and to identify potential factors that may mediate the results for RQ1 and RQ2. However, as these results are not the primary focus of our study, we detail them in Appendix A.1.
@@ 3.5. Analysis
When designing the experiments, we considered the effects of agent(s) on the participants based on different topics and whether the agents agreed or disagreed with the participants. Therefore, we examined results in each of the following scenarios: Topic 1, Topic 2, Same Stance, and Different Stance.
To verify the assumption required for parametric analyses, we first performed a Levene test and found that all items met the homogeneity of variances. We then conducted a Shapiro-Wilk test to assess data normality. For measures that met the normality assumption, we used one-way ANOVA followed by Tukey’s test for post-hoc analysis. For measures that did not meet the normality assumption, we applied the Kruskal-Wallis test and performed Dunn’s test as a post-hoc analysis.
To evaluate opinion change, we conducted two analyses: (1) examining whether users’ opinions shifted towards the bots’ stance, and (2) assessing changes in the polarization level of users’ opinions. These two analyses are grounded in our hypotheses: we expected that in cases where the agents disagreed with the user, the multi-agent condition would lead to greater opinion change toward the agents’ stance; whereas when the agents agreed with the user, the presence of multiple agents would reinforce the user’s views and lead to increased opinion polarization. These hypotheses are described in Section 2.3.
@@ 3.5.1. Opinion Change Analysis
For the first analysis, we calculated opinion change based on the stance of the bots, where a higher value of ”Opinion Change” indicates a greater shift of participants’ opinions towards the bots’ stance. If the bot supported the topic, we expected participants’ ratings to increase after the conversation; conversely, if the bot opposed the topic, we anticipated a decrease. Thus, we defined opinion change, $\Delta O$ as follows:
[…]
where $O_{\text{post}}$ and $O_{\text{pre}}$ represent participants’ topic ratings in the post-survey and pre-survey, respectively. This approach captures the direction and magnitude of participants’ opinion changes toward the agent(s).
3.5.2. Opinion Polarization Analysis
@@ 4. Results
Before analyzing the variables related to our RQs, we conducted an analysis of the control variables and found no significant differences between groups (as shown in Appendix A.2). As a result, we did not include these control variables in the subsequent analyses.
@@ 4.1.1. Quantitative
Opinion Change. We found significant results in opinion change towards agents, as shown in Figure 6. The differences were observed under Topic 2 (Kruskal-Wallis: H(2)=7.757, p¡0.05), and also when the agents hold different stances towards the users (Kruskal-Wallis: H(2)=6.937, p¡0.05). Under Topic 1 (Kruskal-Wallis: H(2)=0.174, p=0.916) and when the agents holding same stances with the users (One-way ANOVA: F(2, 91) = 0.620, p = 0.54), no significant difference was observed.
@@
By conducting post-hoc analysis, we found that, for Topic 2, participants in the 3-agent group shifted their opinions more towards the bots’ (M=0.327, SD=0.751) than those in the 1-agent group (M=-0.219, SD=0.716; Dunn’s Test: p¡0.05). Similarly, when the bots disagreed with the participants, there was a greater shift in the 3-agent group (M=0.509, SD=1.021) than in the 1-agent group (M=-0.161, SD=0.802; Dunn’s Test: p¡0.05). These results indicate that, after conversations with agent(s), users changed opinions more towards the bots’ stance when discussing with 3-agents than a single agent.
@@
Opinion Polarization. We observed a significant difference in the polarization of participants’ opinions when the agents had the same opinion as them (Kruskal-Wallis: H(2) = 9.962, p¡0.01), as shown in Figure 7. The other results were as follows: Topic 1 (Kruskal-Wallis: H(2) = 4.434, p = 0.108), Topic 2 (F[2,91]=1.336, p=0.267), and when agents disagreed with the participants (Kruskal-Wallis: H(2) = 1.899, p = 0.386).
@@
In a post-hoc analysis, we observed that when the agents and participants agreed, the 5-agent group participants became more polarized (M=0.580, SD=0.576, Dunn’s test: p¡0.01) than the 3-agent participants (M=0.132, SD=0.640). The 1-agent group (M=0.252, SD=0.383) did not differ significantly from either the 3-agent (p=0.134) or 5-agent groups (p=0.096). This indicates that as the number of agents increases from three to five, the polarization level of the participants increases significantly. This finding supports RQ1, suggesting that multiple agents can indeed create greater opinion change.
In contrast to the ”agree” condition, no significant changes were observed when the agents and participants disagreed. This was to be expected, as when the agents held the same stance as a user, they would naturally be more likely to nudge the user’s stance further in its existing direction rather than against it (i.e., more strongly agreeing or disagreeing with the topic), causing greater polarization. In contrast, when the agents held a different stance to the user, they were more likely to bring the user back towards the neutral point (e.g., causing a disagreeing user to disagree less). Thus, this did not increase polarization.
@@
Figure 6. Opinion change after agent interaction. Each boxplot shows the distribution of participants’ opinion changes across three agent conditions (1 Agent, 3 Agents, 5 Agents). The yellow circle represents the mean, and the horizontal gray dashed line indicates the overall median. Participants in the 3-agent group had a significantly greater opinion shift towards the agents’ opinions than those in the 1-agent group, when discussing Topic 2 and when the agents and the user had different stances.
@@
Figure 7. Opinion polarization after agent interaction. Each boxplot shows the distribution of participants’ opinion polarization level changes across three agent conditions (1 Agent, 3 Agents, 5 Agents). The yellow circle represents the mean, and the horizontal gray dashed line indicates the overall median. Participants in the 5-agent group showed significantly greater polarization in their opinions than those in the 3-agent group, when the agents and the user had the same stances.
@@ 4.1.2. Qualitative
The qualitative data analysis aimed to address the question: How do participants describe changes in their opinions as influenced by the agents? This analysis was conducted to complement and validate the quantitative results from the previous section.
To uncover participants’ understanding of whether their opinions had changed, we analyzed responses to the open-ended question, ”In what ways did agent(s) influence your opinion during the discussion?”, by thematic analysis (Braun and Clarke, 2006), specifically in the cases where the agents disagreed with the user. Participants who responded that they were ”not persuaded” were coded as ”no opinion change”, while those who mentioned agreeing with the agents or changing their minds to some extent were coded ”some opinion change”.
@@ 4.2.1. Quantitative
There are two types of social influence measured: informational influence and normative influence. We will report on them accordingly.
@@
Normative Influence. For normative influence, we found significant differences across both topics (Topic 1: H(2)=6.571, p¡0.05; Topic 2: H(2)=9.111, p¡0.05), when the bots’ stances were the same as the users (Same stance: H(2)=8.708, p¡0.05) and were specifically notable when the bots’ stances differed from those of the users (Different stance: H(2)=10.100, p¡0.01). These results are presented in Figure 9.
The post-hoc analysis revealed that the 5-agent group participants felt significantly more normative influence than those in the 1-agent group across all four scenarios. For Topic 1, participants felt significantly more normative influence in the 5-agent (M=3.833, SD=0.854; Dunn’s test: p¡0.05) group as compared to those in the 1-agent group (M=3.129, SD=1.162). A similar pattern emerged for Topic 2, where significantly more normative influence was found in the 5-agent (M=3.717, SD=1.179; Dunn’s test: p¡0.01) as compared to the 1-agent group (M=2.919, SD=1.081). When the bot had the same stance as the user, significantly more normative influence was observed with 5 agents (M=3.750, SD=1.015; p¡0.05) as compared to with 1 agent (M=2.935, SD=1.138). Only when the bot had a different stance from the user, significantly more normative influence was observed with 3 agents (M=3.848, SD=0.906; p¡0.05) and 5 agents (M=3.750, SD=1.015; p¡0.05) as compared to with 1 agent (M=2.935, SD=1.138).
@@
Figure 8. Informational influence perceived by participants. Each boxplot shows the distribution of participants’ perceived informational influence across three agent conditions (1 Agent, 3 Agents, 5 Agents). The red circle represents the mean, and the horizontal gray dashed line indicates the overall median. In all conditions, no significant differences in informational influence were found.
Figure 9. Normative influence perceived by participants. Each boxplot shows the distribution of participants’ perceived normative influence across three agent conditions (1 Agent, 3 Agents, 5 Agents). The red circle represents the mean, and the horizontal gray dashed line indicates the overall median. Participants in the 5-agent group felt significantly more normative influence than those in the 1-agent group in all four conditions. Additionally, participants in the 3-agent group also felt significantly more influenced than those in the 1-agent group when the agents and the user had different stances.
@@ 4.2.2. Qualitative
To explore how agents affect users’ perceived social influence and opinion change, we analyzed users’ responses to open-ended questions related to 1) describing impressions of the agent(s) and 2) describing reasons behind their opinion changes, to uncover the underlying mechanism of social influence. We identified two potential factors contributing to users’ feelings of being influenced by the agent(s).
@@
Firstly, participants perceived the accuracy of the agents’ arguments as a key factor in being influenced. Across all three conditions, they frequently described the arguments as accurate, well-informed, and persuasive. This credibility encouraged participants to reconsider their views. For example, P67 from the 5-agent condition noted, ”I was persuaded by their insightful thoughts. They offered solid arguments about security, stress-free experiences, low accident rates, and more.” To evaluate how this observation is different across three conditions, we conducted a similar round of analysis as in Section 4.1.2 to understand how users perceived the arguments presented by the agents. We analyzed responses to the question, ”Do you think the arguments presented by [the agents] are accurate and convincing, and why?” We deductively coded each response as ”yes”, ”no”, or ”unsure”, with the last code being for responses that were ambiguous in nature (e.g. ”they were accurate but not convincing”). For topics where the agents’ stances agreed with the participants’, many participants considered the arguments convincing: 18/31 (58%) in the 1-agent scenario, 18/33 (54%) with 3 agents, and 15/30 (50%) with 5 agents. For topics where the agents and participants disagreed, some participants still found the arguments convincing (10/31 (32%) in the 1-agent scenario, 9/33 (27%) with 3 agents, and 11/30 (36%) with 5 agents), with no noticeable trend across the three conditions.
@@
Secondly, we observed that participants perceived group-induced social pressure as a critical factor to change their opinions as the number of agents increased. We analyzed the responses to the question, ”During the conversation, did you feel any pressure to agree with the agent(s), and why?” Responses that mentioned feeling pressure or otherwise alluding to it were coded ”yes”, while those reporting no pressure were coded ”no”. Under this coding scheme, in the 1-agent group, only 1 participant (3%) reported feeling pressure; this rose to 4 participants (12%) with 3 agents and 6 participants (20%) with 5 agents.
@@
We further analyzed the responses coded ”yes” to determine the exact reasons for this sense of pressure, and found a desire for affiliation from multi-agent conditions, where some participants reported feelings of social exclusion during discussions with the agents. For example, when asked whether they felt pressure to agree with the agents, P51 from the 3-agent condition said, ”they say at the end something similar to ’Bella, Cody and Nathan all agree’ which for a split second made me feel like an outsider.” P57 from the 3-agent condition also shared, ”The host made it a point that the other agents agreed with each other while I was kind of the odd one out.” Similarly, P77 from the 5-agent condition expressed feelings of exclusion, stating, ”It felt like they were in a high school clique that I wasn’t a part of, kinda like mean girls but with AI.” This perception of social exclusion was almost exclusive to scenarios where the agents disagreed with users.
@@ 4.3. RQ3: Which user demographics are more likely to be influenced by multi-agent systems?
The results showed some significant relationships between IVs and normative influence and opinion change, but no significant effects for opinion polarization. Therefore, we report only the findings on normative influence and opinion change here. Additionally, since no significant relationships were found between ”gender” and the other dependent variables, we will focus exclusively on the findings related to ”age” and ”education.”
@@ 4.3.1. Age
When analyzing the opinion change, the model showed a slight negative association between age and opinion change in 3-agent condition (Topic 2: $\beta$ = -0.0263, $p$ ¡ 0.05, $R^{2}$ = 0.189; Different Stance: $\beta$ = -0.0285, $p$ = 0.053, $R^{2}$ = 0.119), while no trend was observed in 1-agent condition and 5-agent condition. Although the effect was not statistically significant at the conventional 0.05 level, the p-value suggests a marginal trend. This indicates that younger participants tended to change their opinion more compared to older participants in the 3-agent condition. Similarly, the linear regression analysis showed no significant trend in the 1-agent or 3-agent conditions between age and normative influence. However, a negative association was observed between age and normative influence in the 5-agent condition (Topic 1: $\beta$ = -0.0347, $p$ ¡ 0.05, $R^{2}$ = 0.148; Different Stance: $\beta$ = -0.037, $p$ = 0.059, $R^{2}$ = 0.121). The coefficient indicates that younger participants tended to report higher normative influence compared to older participants in the 5-agent condition.